OpenStudy (anonymous):

How many sides does a polygon with 44 right angles have?

OpenStudy (anonymous):

n^2 - 3n / 2 = 44 n^2 - 3n = 88 n^2 - 3n - 88 = 0 (n - 11) (n + 8) = 0 n = 11 or -8 We can ignore the -8 term as we can't have a negative number of sides, so it must be 11, therefore your polygon has 11 sides for 44 diagonals.